# DiffuGR: Generative Document Retrieval with Diffusion Language Models

### v4

@@ 3.2.2. Linguistic DocIDs
With the two simple methods mentioned above, we can easily construct DocIDs for DiffuGR. Note that the focus of this paper lies in non-autoregressive generation of DocIDs other than the construction. Therefore, using such simple DocID construction methods allows us to better validate the effectiveness of diffusion-based generation. As for how to construct more suitable DocIDs for diffusion models, we will explore this problem in future works.
@@ 3.3. Diffusion for Generative Retrieval
To generate DocIDs for a given query, DiffuGR employs masked diffusion language models, which differ from standard auto-regressive models. Specifically, instead of predicting the next token sequentially, the masked diffusion language model learns to iteratively denoise partially masked sequences until a complete identifier sequence is obtained. Masked diffusion language models introduce a forward process that gradually adds noise to the sequence and learn a corresponding reverse process to generate DocIDs.
@@ 3.3.1. Forward process
Let $V$ and $L$ denote the vocabulary size of a masked diffusion language model and the length of input sequence, respectively. Given a sequence $\bm{x_{0}}\in\{0,1,\dots,V-1\}^{L}$ and a noise level $t\in[0,1]$, the forward process randomly and independently masks the tokens in the sequence, formulated as follows:
(4) $q_{t|0}(\bm{x_{t}}|\bm{x_{0}})$ $=\prod_{i=0}^{L-1}q_{t|0}(\bm{x_{t}^{i}}|\bm{x_{0}^{i}})$
@@
where $\bm{x^{i}}$ denotes the $i$-th element of $\bm{x}$, $m$ denotes the mask token (devlin2018bert), $\bm{x_{t}}$ denotes the noisy data at time $t$ and $q_{0}(\cdot)$ is the data distribution $p_{\textrm{data}}(\cdot)$.
@@ 3.3.2. Reverse process
Here $q_{0|t}(\cdot|\cdot)$ is the data prediction model to be learned. Notably, ou2024your revealed an intrinsic property that $q_{0|t}(\cdot|\cdot)$ can be represented by conditional distributions on clean data $p_{\text{data}}(\cdot|\cdot)$ independently from $t$. Formally,
@@
where $\bm{x_{t}}^{\text{UM}}$ collects all unmasked tokens in noisy data $\bm{x_{t}}$ and $p_{\text{data}}(\cdot|\cdot)$ is irrelevant to tt. [^2] For example, if $\bm{x_{t}}=[1,3,m,2]$, then $\bm{x_{t}}^{\text{UM}}=[1,3,\cdot,2]$ and $p_{\text{data}}(\cdot|[1,3,\cdot,2])$ is irrelevant to $t$.
@@ 3.3.3. Training Objective
(9) $\begin{split}-\log&p_{\bm{\theta}}(\bm{x_{0}})\leq\\ &\int_{0}^{1}\frac{1}{t}\mathbb{E}_{q(\bm{x_{t}}|\bm{x_{0}})}\left[\sum_{\{i|\bm{x_{t}^{i}}=m\}}-\log p_{\bm{\theta}}(\bm{x_{0}^{i}}|\bm{x_{t}})\right]dt\triangleq\mathcal{L}.\end{split}$
To put it simply, given a sequence $\bm{x_{0}}$, we randomly mask a subset of tokens according to a corruption process $q_{t|0}(\bm{x_{t}}|\bm{x_{0}})$, where $\bm{x_{t}}$ denotes the corrupted sequence at time $t$. This objective resembles denoising autoencoding, but the iterative refinement aligns naturally with the diffusion process, where the model progressively recovers signal from noise. As for generative retrieval task, the training data contains query-DocID pairs $(q,\mathbf{z})$ and document-DocID pairs $(d,\mathbf{z})$. We only randomly mask the DocID sequences $\mathbf{z}$ for training the masked diffusion language model. During training, both categories of DocIDs are used to construct supervision pairs $(q,\mathbf{z})$. Queries are paired with their corresponding document identifiers, and the model learns to generate $\mathbf{z}$ from $q$.
@@ 3.4. Model Inference
At inference time, the query is concatenated with fully masked DocID placeholders, and the model attempts to generate DocID tokens in parallel and refine the generation through diffusion-based denoising. The total number of denoising iteration is a hyperparameter, which naturally provides diffusion language models with a trade-off between efficiency and quality. Following nie2025large; ye2025dream, we employ uniformly distributed $t$ by default. Due to the inherent properties of the masked diffusion language model, the generation length $l$ is also a predefined hyperparameter. In our work, we set $l$ to the maximum length of the DocIDs.
@@ 3.4.1. Diffusion Denoising Strategies
At the first step, we feed both $q$ and $\mathbf{z}^{1}$, where the superscript of $\mathbf{z}^{1}$ means that the DocID tokens are all masked. Then, for steps from time $t\in(0,1]$ to $s\in[0,t)$, we feed both $q$ and $\mathbf{z}^{s}$ into the masked diffusion language model. Here, $\mathbf{z}^{s}$ denotes the DocID tokens in timestamp $s$, where some tokens remain masked. The model then predicts all currently masked DocID tokens simultaneously. Subsequently, DiffuGR re-masks a fraction of the newly predicted tokens, with an expected proportion of s/t{s}/{t}. This ensures that the transition in the reverse process is consistent with the forward process, enabling accurate sampling.
When it comes to specific implementation of re-masking, a variety of different strategies can be used. In this work, DiffuGR involves four denoising scheduling methods that determine which token to be re-masked and regenerated during the reverse diffusion process:
(1) Random. The random decoding strategy follows a purely random generation order (austin2021structured), which serves as the default baseline for diffusion-based language models. This strategy introduces maximal stochasticity into the denoising trajectory but may lead to suboptimal recovery in structured tasks such as GR.
(2) Maskgit plus. This method adopts the confidence-based token scheduling from (chang2022maskgit), where tokens with the lowest predicted confidence are re-masked. This deterministic progression encourages stable convergence and reduces accumulation of early-stage errors.
(3) Top-$k$ margin. This method extends the confidence-based scheduling by ranking tokens according to margin confidence (i.e., the difference between top-1 and top-2 probabilities) (kim2025train). This method emphasizes the relative certainty of token predictions, prioritizing positions with more decisive model beliefs for generation.
(4) Entropy This method computes the entropy of the token probability distribution and re-masks tokens with the highest entropy, effectively denoising regions where the model exhibits higher confidence firstly. Compared to the margin-based strategy, entropy-based scheduling provides a smoother and information-theoretic measure for generation confidence.
Compared to auto-regressive decoding, such above paradigm offers more flexibility in parallel token updates, and is well-suited for generating structured DocIDs with multiple discrete components. At inference time, given a query, the model outputs a DocID, which is then mapped to the corresponding document in $\mathcal{D}$. In Section 5.4.1, experiments were conducted to investigate the influence of different denoising strategies.
@@ 3.4.2. Pseudo Beam Search for DiffuGR
While DiffuGR offers efficient parallel token generation, it inherently lacks the sequential tree structure that enables candidate exploration in beam search. Consequently,DiffuGR cannot directly leverage beam expansion to produce multiple candidate DocIDs. To address this limitation, we propose two complementary strategies that approximate the effect of beam search within the diffusion-based framework while maintaining its parallel decoding nature.
The first strategy leverages query augmentation, where a single query is reformulated into multiple semantically equivalent but lexically diverse variants through paraphrasing models. These augmented queries are jointly fed into the diffusion generator as independent inputs. Each variant undergoes the same denoising trajectory, yielding a set of candidate DocIDs that capture multiple plausible retrieval hypotheses. This procedure effectively mimics beam search exploration by increasing input diversity rather than branching during generation. In this paper, DeepSeek-V3-0324 is used to conduct the query augmentation.
The second strategy exploits intermediate denoising states within the diffusion process itself. Given a total of $T$ denoising steps, the model produces $T$ intermediate DocID candidates corresponding to different noise levels. By retaining and evaluating these intermediate outputs, DiffuGR can explore multiple decoding hypotheses at negligible additional cost. However, these intermediate DocIDs may contain duplicates or invalid tokens due to stochastic denoising noise, necessitating a lightweight post-filtering mechanism.
The above techniques enable DiffuGR to approximate candidate diversity of beam search without sacrificing its parallel generation property, thus improving retrieval robustness and recall metrics.
@@ 4. Experimental Setup
In this section, we describe the detail of experimental settings.
4.1. Datasets
We conduct experiments on two well-known document retrieval datasets: NQ (Kwiatkowski2019NaturalQA) and MS MARCO (Campos2016MSMA). Statistics of these datasets are summarized in Table 1.
Table 1. Statistics of datasets.
Dataset # Docs # Test queries # Train pairs
NQ320K 109,739 7,830 307,373
MS MARCO 323,569 5,187 366,235
NQ320K. NQ320K is a popular dataset for evaluating generative retrieval models (Tay2022TransformerMA; Wang2022ANC). It is based on the Natural Questions (NQ) dataset proposed by Google (Kwiatkowski2019NaturalQA). NQ320k consists of 320k query-document pairs, where the documents are gathered from Wikipedia pages, and the queries are natural language questions.
MS MARCO. MS MARCO is a collection of queries and web pages from Bing search. Similarly to NQ320k (Zhou2022UltronAU), we sample a subset of labeled documents and use their corresponding queries for training. We evaluate the models on the queries of the MS MARCO dev set and retrieval on the sampled document subset.
4.2. Baselines and Evaluation Metrics
Three types of baselines are considered in our work, including sparse retrieval methods, dense retrieval methods, and generative retrieval methods. More details about baselines can be found in Appendix A.1. On NQ320K, we use Recall@{1,5,10} and Mean Reciprocal Rank (MRR)@10 as evaluation metrics, following (mekonnen2025lightweight). On MS MARCO, we use Recall@{1,5,10} and MRR@10.
@@ 4.3. Implementation Details
Models and Hyperparameters. In our experiments, we utilize LLaDA (nie2025large) and Dream (ye2025dream) as the diffusion language models. As for the learnable DocIDs, the length of DocIDs is set as $l=3$ and the size of $i$-th codebook is $K_{i}=512,i\in\{1,2,3\}$. The maximum number of tokens is $12$ for linguistic DocIDs. We optimize the model using AdamW and set the learning rate to 5e45e-4. The batch size is $32$, and the total number of training epochs is $30$.
Data Augmentation. Following previous work (Zhuang2022BridgingTG; Wang2022ANC; Wang2021GPLGP), we use query generation models to generate synthetic (query, document) pairs for training data augmentation. Following sun2023learning, we use the pre-trained query generation model from DocT5Query (Cheriton2019FromDT) to augment the NQ320K and MS MARCO datasets. In query generation, we use nucleus sampling with parameters $p=0.8,t=0.8$ and generate ten queries for each document in the collection.
5. Results and Analysis
In this section, we present the experimental results to systematically evaluate the performance of DiffuGR and analyze the contribution of its core components. Our evaluation aims to address the following key research questions:
  1. Q1
    How does DiffuGR perform compared to baseline methods?
  2. Q2
    How does DiffuGR balance the quality–latency tradeoff?
  3. Q3
    How does DiffuGR generate top-$k$ results without beam search?
  4. Q4
5.1. Performance Comparison (Q1)
To answer Q1, we compare DiffuGR against various baselines across sparse, dense, and generative retrieval paradigms.
Table 2. Performance comparison on NQ320K. The best results are shown in bold. The second-best values are underlined. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. Abbreviations denote DocIDs used for GR baselines: SI – Semantic ID; PQ – Product Quantization; NG – N-grams; TU – Title + URL. Results of baselines are copied from  (mekonnen2025lightweight). R@$k$ is short for Recall@$k$.
@@ 5.1. Performance Comparison (Q1)
Sparse retrieval
@@
DDRO (TU) 40.86 53.12 55.98 45.99
DDRO (PQ) 48.92 64.10 67.31 55.51
Ours
LLaDA-Learnable 62.63 65.97 66.01 64.20
LLaDA-Linguistic 66.15 66.83 66.90 66.26
Dream-Learnable 65.63 68.23 68.54 66.62
Dream-Linguistic 69.47 69.61 69.78 69.57
Results on NQ320K. Table 2 summarizes the performance comparison on the NQ320K dataset. We draw the following observations:
(1) The proposed DiffuGR (Dream-Linguistic) significantly outperforms all generative retrieval baselines, achieving 20.55% and 14.06% higher R@1 and MRR@10, respectively, compared to the best-performing baseline, DDRO (mekonnen2025lightweight). This highlights the effectiveness of the diffusion process in accurate single-document identification.
(2) Different types of DocID affect the performance of DiffuGR. Overall, using linguistic DocIDs yields better performance than using learnable DocIDs. This is likely because linguistic DocIDs can directly leverage the semantic space of existing models, whereas learnable DocID require vocabulary expansion and semantic remapping, which further processes learning difficulty.
(3) Compared with R@1, our method achieves marginal improvement on R@5 and R@10 metrics. This is attributed to the fact that DiffuGR does not support beam search for top-$k$ generation. We deeply investigate this issue in Section 5.3.
Table 3. Performance comparison on MS MARCO. The best results are shown in bold. The second-best values are underlined. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. Abbreviations denote DocIDs used for GR baselines: SI – Semantic ID; PQ – Product Quantization; NG – N-grams; TU – Title + URL. Results of baselines are copied from  (mekonnen2025lightweight). R@$k$ is short for Recall@$k$.
@@
Sparse retrieval
@@
ROGER-NCI 30.61 59.02 68.78 42.02
ROGER-Ultron 33.07 63.93 75.13 46.35

### v5

@@ 3.2.2. Text-based DocIDs
Here, number-based DocID, e.g, a single number denotes a document, is not used. Although it offers high retrieval efficiency, this type of DocID often exhibits poor generalization performance (Zhang et al., 2025).
@@ 3.3. Diffusion for Generative Retrieval
DiffuGR employs masked diffusion language models to generate DocIDs for a given query. It defines a forward process that progressively masks DocID tokens and learns a reverse denoising process that iteratively reconstructs the complete identifier sequence.
@@ 3.3.1. Forward process
Let $V$ denote the vocabulary size of the masked diffusion language model and $L$ the length of the input sequence. Given a sequence $\bm{x_{0}}\in\{0,1,\dots,V-1\}^{L}$ and a noise level $t\in[0,1]$, the forward process randomly and independently masks the tokens, formulated as follows:
(4) $q_{t|0}(\bm{x_{t}}|\bm{x_{0}})$ $=\prod_{i=0}^{L-1}q_{t|0}(\bm{x_{t}^{i}}|\bm{x_{0}^{i}}),$
@@
where $q_{t|0}(\bm{x_{t}}|\bm{x_{0}})$ denotes the transition probability from the original data $\bm{x_{0}}$ to the noisy data $\bm{x_{t}}$ at time $t$, $\bm{x^{i}}$ denotes the $i$-th element of $\bm{x}$, $m$ denotes the mask token (Devlin, 2018) and $q_{0}(\cdot)$ is the data distribution $p_{\textrm{data}}(\cdot)$.
@@ 3.3.2. Reverse process
Here $q_{0|t}(\cdot|\cdot)$ is the DiffuGR to be learned. Notably, Ou et al. (2024) revealed an intrinsic property of mask diffusion language model that $q_{0|t}(\cdot|\cdot)$ can be represented by conditional distributions on clean data $p_{\text{data}}(\cdot|\cdot)$ independently from $t$, which means:
@@
where $\bm{x_{t}}^{\text{UM}}$ collects all unmasked tokens in noisy data $\bm{x_{t}}$ and $p_{\text{data}}(\cdot|\cdot)$ is irrelevant to $t$. For example, if $\bm{x_{t}}=[1,3,m,2]$, then $\bm{x_{t}}^{\text{UM}}=[1,3,\cdot,2]$ and $p_{\text{data}}(\cdot|[1,3,\cdot,2])$ is irrelevant to $t$.
@@ 3.3.3. Training Objective
(9) $\begin{split}-\log&p_{\bm{\theta}}(\bm{x_{0}})\leq\\ &\int_{0}^{1}\frac{1}{t}\mathbb{E}_{q(\bm{x_{t}}|\bm{x_{0}})}\left[\sum_{\{i|\bm{x_{t}^{i}}=m\}}-\log p_{\bm{\theta}}(\bm{x_{0}^{i}}|\bm{x_{t}})\right]dt\triangleq\bm{L}.\end{split}$
To put it simply, given a sequence $\bm{x_{0}}$, we randomly mask a subset of tokens according to a corruption process $q_{t|0}(\bm{x_{t}}|\bm{x_{0}})$, where $\bm{x_{t}}$ denotes the corrupted sequence at time $t$. This objective resembles denoising autoencoding, but the iterative refinement aligns naturally with the diffusion process, where the model progressively recovers signal from noise. As for generative retrieval task, the training data contains query-DocID pairs $(q,\bm{z})$ and document-DocID pairs $(d,\bm{z})$. We only randomly mask the DocID sequences $\bm{z}$ for training the masked diffusion language model. Queries are paired with their corresponding document identifiers, and the model learns to generate $\bm{z}$ from $q$.
@@ 3.4. Model Inference
At inference time, the query is concatenated with fully masked DocID placeholders, and the model attempts to generate DocID tokens in parallel and refine the generation through diffusion-based denoising. The total number of denoising iteration is a hyperparameter, which naturally provides diffusion language models with a trade-off between efficiency and quality. Following Nie et al. (2025); Ye et al. (2025), we employ uniformly distributed $t$ by default. Due to the inherent properties of the masked diffusion language model, the generation length $l$ is also a predefined hyperparameter. In our work, we set $l$ to the maximum length of the DocIDs.
@@ 3.4.1. Diffusion Denoising Strategies
At the first step, we feed both $q$ and $\bm{z}^{1}$, where the superscript of $\bm{z}^{1}$ means that the DocID tokens are all masked. Then, for steps from time $t\in(0,1]$ to $s\in[0,t)$, we feed both $q$ and $\bm{z}^{s}$ into the masked diffusion language model. Here, $\bm{z}^{s}$ denotes the DocID tokens in timestamp $s$, where some tokens remain masked. The model then predicts all currently masked DocID tokens simultaneously. Subsequently, DiffuGR re-masks a fraction of the newly predicted tokens, with an expected proportion of s/t{s}/{t}. This ensures that the transition in the reverse process is consistent with the forward process, enabling accurate sampling.
When it comes to specific implementation of re-masking, a variety of different strategies can be used. In this work, DiffuGR explores four denoising scheduling methods that determine which token to be re-masked and regenerated during the reverse diffusion process: (1) Random. The random decoding strategy follows a purely random generation order (Austin et al., 2021), which serves as the default baseline for diffusion-based language models. (2) Maskgit plus. This method adopts the confidence-based token scheduling from (Chang et al., 2022), where tokens with the lowest predicted confidence are re-masked. (3) Top-$k$ margin. This method extends the confidence-based scheduling by ranking tokens according to margin confidence (i.e., the difference between top-1 and top-2 probabilities) (Kim et al., 2025). (4) Entropy. Tokens with high entropy, indicating low model confidence, are re-masked and re-predicted in later diffusion iterations, whereas low-entropy tokens are preserved to stabilize the denoising process. In Section 5.4.1, experiments were conducted to investigate the effects of different denoising strategies on the model performance.
@@ 3.4.2. Pseudo Beam Search for DiffuGR
While DiffuGR offers efficient parallel token generation, it inherently lacks the fixed left-to-right generation order that enables candidate exploration in beam search. Consequently, DiffuGR cannot directly leverage standard beam search to generate multiple candidate DocIDs. To address this limitation, we explored three strategies to approximate the execution of beam search: (1) The first strategy leverages query augmentation, where a single query is reformulated into multiple semantically equivalent but lexically diverse variants through a language model. (2) Candidate Trajectory Extraction (CTE) is the second strategy, which exploits intermediate denoising states within the diffusion process itself. Given a total of $T$ denoising steps, the model produces $T$ intermediate DocID candidates corresponding to different noise levels. By retaining and evaluating these intermediate outputs, DiffuGR can explore multiple decoding hypotheses at negligible additional cost. (3) The third strategy is an adapted beam search for diffusion model (Shi et al., 2025), which has been exploited in generative recommendation. At each iteration, it first identify the position associated with the highest model confidence for generation, followed by conducting beam expansion at this selected position. Through successive rounds of this iterative process, the final recommended top items $B$ are obtained.
The above techniques enable DiffuGR to approximate candidate diversity of beam search without sacrificing its parallel generation property, thus improving retrieval robustness and recall metrics. However, due to the incompatibility with constrained decoding, None of these methods can guarantee that the generated DocIDs are valid and non-duplicate. To mitigate these issues, DiffuGR performs validity and deduplication checks on the generated DocIDs. For each invalid DocID, DiffuGR searches for the nearest valid DocID based on Hamming distance.
@@ 4. Experimental Setup
4.1. Datasets and Evaluation Metrics
We conduct experiments on two well-known document retrieval datasets: NQ320K (Kwiatkowski et al., 2019) and MS MARCO (Campos et al., 2016).
NQ320K. NQ320K is a widely applied benchmark for evaluating retrieval models (Tay et al., 2022; Wang et al., 2022b), based on the Natural Questions (NQ) dataset (Kwiatkowski et al., 2019). NQ320k consists of 320k query-document pairs, where the documents are gathered from Wikipedia pages, and the queries are natural language questions. Note that our study follows the data processing pipeline as described in DDRO (Mekonnen et al., 2025), and obtained approximately 100k documents.
MS MARCO. MS MARCO is a collection of queries and web pages from Bing search. Similarly to NQ320k, we sample a subset of labeled documents and use their corresponding queries for training. We evaluate the models on the queries of the MS MARCO dev set and retrieval on the sampled document subset.
Evaluation Metrics. On NQ320K and MS MARCO, we use R@{1,5,10} and MRR@10 as evaluation metrics, following (Mekonnen et al., 2025). R@100 and nDCG@10 are also used in some analytical experiments.
4.2. Baselines
Three types of baselines are considered in our work, including sparse retrieval methods, dense retrieval methods, and generative retrieval methods.
The sparse retrieval baselines are as follows: BM25 uses the tf-idf feature to measure term weights and we use the implementation from http://pyserini.io/. DocT5Query expands a document with possible queries predicted by a finetuned T5 with this document as the input.
The dense retrieval baselines are as follows: DPR (Karpukhin et al., 2020), a dual-encoder model using the representation of the [CLS] token of BERT. ANCE (Xiong et al., 2021), an asynchronously updated ANN indexer is utilized to mine hard negatives for training a RoBERTa-based dual-encoder model. Sentence-T5 (Ni et al., 2022), a dual-encoder model that uses T5 to produce continuous sentence embeddings. RepBERT (Zhan et al., 2020), a BERT-based model that generates fixed-length contextualized embeddings with query-document relevance computed via inner product similarity.
The generative retrieval baselines are as follows: DSI (Tay et al., 2022), which represents documents using hierarchical $K$-means clustering results, and indexes documents using the first 32 tokens as pseudo-queries. SEAL (Bevilacqua et al., 2022) uses arbitrary n-grams in documents as DocIDs, and retrieves documents under the constraint of a pre-built FM-indexer. DSI-QG (Zhuang et al., 2022) uses a query generation model to augment the document collection. Ultron (Zhou et al., 2022b) uses a three-stage training pipeline and represents documents through three types of identifiers. ROGER (Zhou et al., 2024),which transfers knowledge from a dense retriever to a generative retriever via knowledge distillation. MINDER (Li et al., 2023),which assigns multiple identifiers, including titles, n-grams, and synthetic queries, to documents and pairs them for indexing. LTRGR (Li et al., 2024b),which trains on pairwise relevance objectives using margin-based ranking loss for optimization. GenRRL (Zhou et al., 2023),which incorporates pointwise, pairwise, and listwise relevance optimization through reinforcement learning, using document summaries and URLs as DocIDs. DDRO (Mekonnen et al., 2025), which enhances generative retrieval by directly aligning generation with document-level relevance estimation.
@@ 4.3. Implementation Details
Models and Hyperparameters. In our experiments, we utilize LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025) as the diffusion language models. As for the learnable DocIDs, the length of DocIDs is set as $l=3$ and the size of $i$-th codebook is $K_{i}=512,i\in\{1,2,3\}$. For the MS MARCO dataset, the length of learnable DocIDs is set as $l=4$. Beam size is set to 15. The maximum number of tokens is $12$ for text-based DocIDs. We optimize the model using AdamW and set the learning rate to 5e45e-4. DeepSeek-V3 is used for the query augmentation.
Data Augmentation. Following previous work (Zhuang et al., 2022; Wang et al., 2022b, a), we use query generation models to generate synthetic (query, document) pairs for training data augmentation. Following Sun et al. (2023), we use the pre-trained query generation model from DocT5Query (Cheriton, 2019) to augment the NQ320K and MS MARCO datasets. In query generation, we use nucleus sampling with parameters $p=0.8,t=0.8$ and generate ten queries for each document in the collection.
Table 1. Performance comparison on NQ320K. The best results are shown in bold. The second-best values are underlined. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. Abbreviations denote DocIDs used for GR baselines: SI – Semantic ID; PQ – Product Quantization; NG – N-grams; TU – Title + URL. Some results of baselines are copied from  (Mekonnen et al., 2025), and * denotes that the results are reproduced using public code. R@$k$ is short for Recall@$k$.
@@
Sparse retrieval & Dense retrieval
@@
DDRO (TU)* 39.75 52.07 54.41 44.82
DDRO (PQ)* 48.10 62.58 66.78 54.25
LLaDA-Learnable 62.64 76.05 77.79 66.10
LLaDA-Text 66.29 79.06 81.25 71.86
Dream-Learnable 65.63 78.59 79.70 69.13
Dream-Text 69.73 80.90 82.82 74.65
Table 2. Performance comparison on MS MARCO. The best results are shown in bold. The second-best values are underlined. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. Abbreviations denote DocIDs used for GR baselines: SI – Semantic ID; PQ – Product Quantization; NG – N-grams; TU – Title + URL. Some results of baselines are copied from  (Mekonnen et al., 2025), and * denotes that the results are reproduced using public code. R@$k$ is short for Recall@$k$.
@@
Sparse retrieval & Dense retrieval
@@
ROGER-NCI (SI) 30.61 59.02 68.78 42.02
ROGER-Ultron (TU) 33.07 63.93 75.13 46.35